OpenStudy (anonymous):

A pool in the shape of a rectangular prism has a volume of 2100m^3. The dimensions of the pool are x metres deep by 25x metres long by 10x+1 metres wide. What are the actual dimensions of the pool?

OpenStudy (campbell_st):

ok... you need to set up the equation \[V = l \times w \times h\] and you know V = 2100 so \[2100 = x \times 25x \times(10x + 1)\] simplify the right hand side of the equation and then solve for x

OpenStudy (anonymous):

How do I simplify the right hand side